OpenStudy (anonymous):

find the intercepts and use them to graph the equation 3y - 9 = 3x

OpenStudy (anonymous):

3y=3x+9 y = x + 3 x + 3 = 0 x = -3 (x intercept0 y = 0 + 3 y = 3 (y intercept)